An architect designs a rectangular flower garden such that the width is exactly​ two-thirds of the length. If 420 feet of antique picket fencing are to be used to enclose the​ garden, find the dimensions of the garden.
Answer by rothauserc(4718)

let w be width of the rectangular flower garden and l be its length, then
:
2(2l/3) + 2l = 420
:
4l/3 +2l = 420
:
multiply both sides of = by 3
:
4l +6l = 1260
:
10l = 1260
:
l = 126
:
w = 2(126)/3 = 84
:
***************************************
length is 126 feet and width is 84 feet
:
check the answer
:
2(126) +2(84) = 420
:
252 +168 = 420
:
420 = 420
:
answer checks
****************************************
: